Conrad Playground Supply underwent a restructuring in 2018. The company conducted a thorough internal audit, during which the following facts were discovered. The audit occurred during 2018 before any adjusting entries or closing entries are prepared.

  1. Additional computers were acquired at the beginning of 2016 and added to the company’s office network. The $40,500 cost of the computers was inadvertently recorded as maintenance expense. Computers have five-year useful lives and no material salvage value. This class of equipment is depreciated by the straight-line method.
  2. Two weeks prior to the audit, the company paid $12,500 for assembly tools and recorded the expenditure as office supplies. The error was discovered a week later.
  3. On December 31, 2017, merchandise inventory was understated by $69,000 due to a mistake in the physical inventory count. The company uses the periodic inventory system.
  4. Two years earlier, the company recorded a 3% stock dividend (1,100 common shares, $1 par) as follows:
Retained earnings 1,100
Common stock 1,100


The shares had a market price at the time of $13 per share.

  1. At the end of 2017, the company failed to accrue $86,000 of interest expense that accrued during the last four months of 2017 on bonds payable. The bonds, which were issued at face value, mature in 2022. The following entry was recorded on March 1, 2018, when the semiannual interest was paid, as well as on September 1 of each year:
Interest expense 129,000
Cash 129,000
  1. A three-year liability insurance policy was purchased at the beginning of 2017 for $69,300. The full premium was debited to insurance expense at the time.


Required:
For each error, prepare any journal entry necessary to correct the error as well as any year-end adjusting entry for 2018 related to the situation described. (Ignore income taxes.) (If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)
Answer #1
Transactions General Journal Debit Credit
a(1) Equipment $40,500
Accumulated depreciation ($40,500/5 =$8,100 ×2 = $16,200) $16,200
Retained earnings $24,300
(Being Equipment recorded)
a(2) Depreciation expense $8,100
Accumulated depreciation $8,100
(Being depreciation expense recorded)
b(1) Cash $12,500
Office supplies $12,500
(Being entry reversed which was recorded wrong)
b(2) Tools $12,500
Cash $12,500

(Being correct entry recorded for tools paid)

c(1) Inventory $69,000
Retained earnings $69,000
(Being inventory was understated, now correctly recorded)
c(2) No Journal entry required
d(1) Retained earnings (1100 × $13 =$14,300 - $1100 =$13,200) $13,200
Paid in capital - ecxess of par $13,200
(Being retained earnings recorded)
d(2) No Journal entry required
e(1) Retained earnings $86,000
Interest expense $86,000
(Being interest expense accrued)
e(2) Interest expense $86,000
Interest payable $86,000
(Being interest expense recorded)
f(1) Prepaid insurance ($69,300/3 =$23,100 ×2 = $46,200
Retained earnings $46,200
(Being prepaid insurance recorded)
f(2) Insurance expense $23,100
Prepaid insurance $23,100
Being insurance expense recorded)